OpenStudy (anonymous):

So in class examples today we were given compounds like calcium Bromide and where to come up with formulas for it like CaBr2. I took notes that this is because Ca has +2 charge and Bromide means that is has -1 charge. Can someone please explain to me how I can gather the info from periodic table about what kind of charge something like Ca has? I just wrote it down from class and do not really understand how to find this info on periodic table.

OpenStudy (anonymous):

Another example was Calcium Carbonate which was then written as CaCo3

OpenStudy (cwrw238):

if you look at the periodic table you see that calcium is in group 2 . It has 2 electrons in its outermost shell and is electro -positive ( on the right of the PT) so its ion will have 2 positive charges

OpenStudy (anonymous):

what do you mean electro - positive?

OpenStudy (cwrw238):

as for Br (bromine) its in Group 7: - 7 electrons in outer shell and is electro negative . To get the maximum stable outer shell which contains 8 electrons it tens to grab one electron in a chemical reaction so its its ion has1 negative charge

OpenStudy (cwrw238):

electropositive elements tend to donate electron and for positive charged ion metallic elements like sodium, calcium, iron are electro positive.

OpenStudy (cwrw238):

Its not as simple as that all elements - for example tin ion can be Sn 2+ or Sn 4+ the above is a general rule

OpenStudy (cwrw238):

There is a concept called Oxidation Number which helps you find the formula of chemical compounds which you'll come across later in your studies.

OpenStudy (anonymous):

Think I a starting to get it. Thanks!

OpenStudy (anonymous):

Im looking up the(+) charged elements right now.

OpenStudy (cwrw238):

CaCO3 Ca ion is 2+ C atom has oxidation number 4 3 oxygen atoms is 3*2- = -6 so CO3 has charge of 4-6 = 2- so Ca and CO3 combine to form the neutral CaCO3

OpenStudy (cwrw238):

some elements can have different charges eg iron can be Fe++ or Fe+++

OpenStudy (cwrw238):

these usually belong to the Transition elements

OpenStudy (cwrw238):

Hope this helps gotta go now